OpenStudy (anonymous):

Find the x and y intercepts and graph the equation plotting the intercepts. 6x-7y=-42

OpenStudy (aroub):

If you want to find the x-intercept you have to substitute y by 0 and the same thing for the y-intercept if you want to find it you have to substitute the x by 0

OpenStudy (anonymous):

for all x intercepts, substitute y for 0 (that means it will be on the x axis) So then: 6x-7(0)=-42 6x=-42 x=-7 So at the point (-7,0) we have an x intercept for all y intercepts, substitute x for 0 (it will be on the y axis) So then: 6(0)-7y=-42 -7y=42 y=-6 So at the point (0,-6) we have a y intercept

OpenStudy (anonymous):

thank u a lot